### v2

# THOR: Tool-Integrated Hierarchical Optimization via RL for Mathematical Reasoning

Abstract
Large Language Models (LLMs) have made remarkable progress in mathematical reasoning, but still continue to struggle with high-precision tasks like numerical computation and formal symbolic manipulation. Integrating external tools has emerged as a promising approach to bridge this gap. Despite recent advances, existing methods struggle with three key challenges: constructing tool-integrated reasoning data, performing fine-grained optimization, and enhancing inference. To overcome these limitations, we propose THOR (Tool-Integrated Hierarchical Optimization via RL). First, we introduce TIRGen, a multi-agent actor-critic-based pipeline for constructing high-quality datasets of tool-integrated reasoning paths, aligning with the policy and generalizing well across diverse models. Second, to perform fine-grained hierarchical optimization, we introduce an RL strategy that jointly optimizes for both episode-level problem solving and step-level code generation. This is motivated by our key insight that the success of an intermediate tool call is a strong predictor of the final answer’s correctness. Finally, THOR incorporates a self-correction mechanism that leverages immediate tool feedback to dynamically revise erroneous reasoning paths during inference. Our approach demonstrates strong generalization across diverse models, performing effectively in both reasoning and non-reasoning models. It further achieves state-of-the-art performance for models of a similar scale on multiple mathematical benchmarks, while also delivering consistent improvements on code benchmarks. Our code will be publicly available at https://github.com/JingMog/THOR.
Figure 1: An overview of our method. The left panel depicts the motivation and challenges, the middle highlights our proposed solution with the TIRGen data construction pipeline and the THOR hierarchical RL framework, and the right panel reports experimental results.
1 Introduction
Large Language Models (LLMs) have achieved remarkable progress, increasingly exhibiting human-like capabilities such as thinking, reflection, and self-correction. They have shown significant improvements in mathematical reasoning, code generation, and autonomous agent tasking (jaech2024openai; guo2025deepseek; yang2025qwen3; team2025kimi).
Recent tool-free methods for enhancing LLMs’ mathematical reasoning can be broadly categorized into search-based methods (besta2024graph; zhang2025lessons; hu2025prm) and reinforcement learning (RL) paradigms (yu2025dapo; yue2025vapo). Despite notable progress, these approaches remain constrained by a fundamental weakness of LLMs. As probabilistic next-token predictors, they inherently struggle with high-precision tasks (chen2022program), such as numerical computation, equation solving, symbolic manipulation (pan2025enhancing), and formal proofs (lewkowycz2022solving). In contrast, programmatic reasoning excels in these domains. Therefore, integrating the semantic reasoning capabilities of LLMs with the precise and verifiable execution of external code-based tools provides a promising pathway to overcome these limitations.
Tool-Integrated Reasoning (TIR) has emerged as a powerful paradigm for enhancing LLM reasoning by enabling them to leverage external tools to augment reasoning (gou2023tora; li2025start). Despite considerable efforts, three core challenges remain: constructing TIR data, performing fine-grained optimization, and enhancing inference. (1) For constructing TIR data, a common approach is to synthesize tool-use data by prompting external powerful models (e.g., GPT-4o) (gou2023tora). However, for reasoning models such as DeepSeek-R1, prompting alone often fails to elicit effective tool use (guo2025deepseek; li2025start). While techniques like START (li2025start) explicitly inject code prompts into the thinking process, purely rule-based approaches struggle to identify suitable insertion positions. Therefore, existing TIR data construction methods suffer from style mismatches and poor applicability to reasoning models. (2) For performing fine-grained optimization, current research primarily employs either SFT or RL. SFT-based methods, like Toolformer (schick2023toolformer) and Aimo-2 (moshkov2025aimo), require large-scale, high-quality demonstration data and often suffer from poor generalization. Existing RL methods (mai2025agent; li2025torl; feng2025retool) typically optimize at the episode-level, overlooking fine-grained updates on specific steps. Although RL is a more scalable alternative, it faces severe sparse reward problems, particularly in long reasoning chains. (3) For enhancing inference, existing methods typically interleave tool calls directly with natural language reasoning in a single pass, thereby overlooking the role of immediate tool feedback in reasoning.
To address these challenges, we propose THOR, a tool-integrated framework designed to enhance the reasoning ability of LLMs. (1) For constructing TIR data, in order to efficiently generate policy-aligned TIR data, we propose TIRGen, an actor-critic-based data construction pipeline. The actor is responsible for generating natural language reasoning steps, while the critic evaluates whether steps can be transformed into executable code and interacts with an external executor to refine the reasoning. This iterative process yields a TIR dataset that is naturally aligned with the actor’s policy and broadly applicable across diverse models and tools. (2) For performing fine-grained optimization, we are motivated by the key insight that the success of an intermediate tool call is a strong predictor of the final answer’s correctness. Our experiments later confirm this insight. Based on this, we introduce a hierarchical RL strategy that combines episode-level and step-level optimization. At the episode-level, we directly optimize for the correctness of the final answer. Concurrently, at the step-level, we apply fine-grained optimization to execution failure steps, specifically enhancing the model’s code generation ability. (3) For enhancing inference, we propose a self-correction mechanism that leverages immediate feedback from tools to dynamically revise its CoT during inference. When code invocation fails, it backtracks and explores alternative reasoning paths, thereby significantly enhancing the model’s reasoning robustness and overall performance.
We evaluate our method on diverse challenging and widely-used benchmarks, including MATH500 (hendrycks2021measuring), AIME 2024 & 2025, AMC, Minerva Math (lewkowycz2022solving), and Olympiad Bench (he2024olympiadbench). THOR establishes a new state-of-the-art (SOTA) result among models of comparable size across architectures and scales, while reducing inference overhead. It further improves performance on code generation benchmarks HumanEval, MBPP (liu2023your), and LiveCodeBench, validating the effectiveness and generalizability of our approach.
Our primary contributions are as follows: 1) Tool-Integrated Data Construction Pipeline. We introduce TIRGen, a pipeline for generating TIR data, applicable across diverse models, and better aligned with the preferences of the policy model. 2) Hierarchical Optimization. We propose a hierarchical reinforcement learning approach that combines episode-level and step-level optimization. 3) Self-correction Inference Enhancement. We introduce a self-correction mechanism that leverages immediate tool feedback to revise reasoning steps during inference. 4) Superior Performance and Broad Generalization. Our approach generalizes across reasoning and non-reasoning models, achieving competitive results on mathematical benchmarks and consistent gains on code tasks.
2 Methodology
2.1 Problem Formulation
In the context of tool-integrated reasoning, an LLM solves mathematical problems by interleaving natural language reasoning with tool invocations. Specifically, we formulate an LLM, parameterized by $\theta$, as a policy $\pi_{\theta}$. Given a problem $q$ and a corresponding instruction $I$, this policy $\pi_{\theta}$ autoregressively generates an entire interaction trajectory $\tau$, which is an alternating sequence of thoughts, actions, and observations:
$\tau=(r^{1},a^{1},o^{1},...,r^{t},a^{t},o^{t},...,r^{n-1},a^{n-1},o^{n-1},r^{n}),$ (1)
where $r^{t}$ is a step of natural language reasoning, $a^{t}$ is an action of tool call, $o^{t}$ is the observation returned by the external execution environment after executing action $a^{t}$, and $n$ is the number of reasoning steps. This process is formulated as an iterative think-act-observe loop. The model incorporates the new observation $o^{t}$ into its context to inform the generation of the subsequent thought $r^{t+1}$ and action $a^{t+1}$. This cycle continues until the model produces the final answer within its last thought $r^{n}$, thereby concluding the trajectory. Formally, the likelihood of generating a specific trajectory $\tau$ is factorized as:
$P_{\pi_{\theta}}(\tau\mid q,I)=P_{\pi_{\theta}}(r^{n}\mid q,I,\mathcal{H}^{1:n-1})\prod_{t=1}^{n-1}\underbrace{P_{\pi_{\theta}}(r^{t}\mid q,I,\mathcal{H}^{1:t-1})}_{\text{Thought}}\underbrace{P_{\pi_{\theta}}(a^{t}\mid r^{t},q,I,\mathcal{H}^{1:t-1})}_{\text{Action}},$ (2)
where $\mathcal{H}^{1:t-1}=\left\{r^{1},a^{1},o^{1},\ldots,r^{t-1},a^{t-1},o^{t-1}\right\}$ denotes the history of the previous interactions. Each term is modeled as a product of token-level probabilities.
2.2 TIRGen: TIR Data Generation Pipeline
Figure 2: The TIR data construction pipeline. In this pipeline, the Actor agent generates reasoning steps. The Critic agent identifies tool-executable steps and converts them into tool-augmented reasoning steps. After multi-stage filtering, we obtain the cold start dataset $\mathcal{D}_{SFT}$.
Existing methods for TIR highlight a significant need for high-quality training data. Most approaches rely on simply prompting external large models to synthesize TIR data for non-reasoning models (gou2023tora; li2025torl), but these approaches fail to generalize effectively to reasoning models such as DeepSeek-R1. Although START (li2025start) constructs long-CoT TIR data using a rule-based prompt-hint approach, the resulting trajectories often contain redundant code invocations. Therefore, existing TIR data construction methods face critical shortcomings, including style mismatches between the generated data and policy models, as well as poor scalability to reasoning models. To overcome these limitations, we introduce TIRGen, an automated TIR data synthesis pipeline based on an actor-critic framework, as illustrated in Figure 2.
In this framework, an Actor agent generates a natural language reasoning step with a maximum length of $L_{\text{step}}$. A Critic agent then evaluates this step to identify operations that can be more efficiently solved with code, such as numerical calculations or equation solving. Upon identifying such an operation, the Critic converts it into an executable Python code snippet while preserving the Actor’s original reasoning logic. The Critic then interacts with a code interpreter to obtain a precise execution result, which replaces the original operation and produces a code-augmented reasoning step. The Actor resumes reasoning from this revised step, and the iterative cycle continues until a complete solution is reached, as described in Algorithm 1. This design yields two key advantages:
  • Policy Alignment. Because the Critic observes only isolated reasoning steps without the full problem statement and final answer, the synthesized data directly reflects the Actor’s intrinsic reasoning preferences. This ensures that the data remains in-distribution, thereby mitigating the performance degradation commonly caused by training on out-of-distribution data (gudibande2023false; chen2024self).
  • Reduced Reliance on Large-scale Models. The Actor is responsible for high-level mathematical reasoning, while the Critic requires only basic instruction-following and code-generation skills. This division of labor decomposes the complex task into subproblems that are collaboratively solved by the two agents.
Algorithm 1 TIRGen: TIR Data Generation Pipeline
1:Input: Actor model $\pi_{\text{actor}}$, Critic model $\pi_{\text{critic}}$, Dataset $\mathcal{D}_{q}$, Code interpreter sandbox $S$.
2:Initialize: Raw cold start dataset $\mathcal{D}_{raw}\leftarrow\emptyset$
3:for question $q\in\mathcal{D}_{q}$ do
4:  Initialize trajectory $\tau\leftarrow(q)$
[…]
6:   $r^{t}\sim\pi_{\text{actor}}(\cdot\mid\tau),\>|r^{t}|\leq L_{\text{step}}$ $\triangleright$ Actor $\pi_{\text{actor}}$ generates a reasoning step
[…]
8:     $r^{t}_{\text{logic}}\leftarrow\text{ExtractLogic}_{\pi_{\text{critic}}}(r^{t})$ $\triangleright$ Step 1: Extract the pure reasoning part
9:     $a^{t}\leftarrow\text{ConvertToCode}_{\pi_{\text{critic}}}(r^{t},r^{t}_{\text{logic}})$ $\triangleright$ Step 2: Convert calculation part to code
10:     $o^{t}\leftarrow S(a^{t})$ $\triangleright$ Step 3: Execute code to get observation
11:     $\tau\leftarrow\tau\oplus(r^{t}_{\text{logic}},a^{t},o^{t})$
12:   else
13:     $\tau\leftarrow\tau\oplus(r^{t})$
14:  $\mathcal{D}_{raw}\leftarrow\mathcal{D}_{raw}\cup\{\tau\}$
[…]
16:Return $\mathcal{D}_{SFT}$
[…]
2.3 Hierarchical RL Training Strategy
Figure 3: A hierarchical optimization framework comprising (a) episode-level RL for mathematical problem solving and (b) step-level optimization for code generation. In addition, we introduce (c) a self-correction mechanism for online error correction during inference.
2.3.1 Cold Start
We initialize our model with a cold start procedure (guo2025deepseek), utilizing the dataset $\mathcal{D}_{SFT}$ generated by TIRGen. The primary objective of this stage is to equip the model with the fundamental patterns of tool invocation. This initial training is crucial, as reasoning models often struggle to invoke code effectively before a cold start. To this end, we perform supervised fine-tuning on the base model $\tilde{\pi}_{\theta}$ with the loss function as follows:
$\mathcal{L}_{\text{SFT}}(\theta)=\mathbb{E}_{(q,I,y)\sim\mathcal{D}_{SFT}}\Big[-\sum_{t=1}^{T}\log\tilde{\pi}_{\theta}(y_{t}\mid q,I,y_{0:t-1})\Big],$ (3)
where $T$ is the length of response. In this way, we obtained the policy model $\pi_{\theta}$ after the cold start.
2.3.2 Hierarchical Reinforcement Learning
In the hierarchical optimization phase, we employ RL to refine the model’s policy for invoking tools to solve complex problems. Conventional RL methods perform optimization at the episode-level, relying on a sparse reward signal derived solely from the correctness of the final outcome (mai2025agent; li2025torl; feng2025retool). Our empirical analysis reveals a crucial insight: the success of an intermediate tool call is a strong predictor of the final answer’s correctness. Motivated by this finding, we propose a hierarchical reinforcement learning framework that combines coarse-grained, episode-level optimization to enhance the model’s problem-solving ability with fine-grained, step-level optimization to improve code generation for individual tool calls, as shown in Figure 3. The observation is validated in Appendix LABEL:appendix:statistical_validation.
Episode-level Optimization. At the episode-level optimization, we adopt GRPO (shao2024deepseekmath) and use the correctness of the final answer as the reward to enhance the tool invocation strategy of the model, as shown in Figure 3(a). The complete episode-level objective is:
$\begin{split}\mathcal{L}^{\text{epis}}_{\pi_{\theta}}(\theta)=&\mathbb{E}[q\sim\mathcal{D}_{RL},\{s_{i}\}_{i=1}^{G}\sim\pi_{\theta}(S|q)]\frac{1}{G}\sum_{i=1}^{G}\Big(\frac{1}{\sum_{t=1}^{|s_{i}|}I(s_{i,t})}\sum_{t:I(s_{i,t})=1}^{|s_{i}|}\\ &\min\big(\frac{\pi_{\theta}(s_{i}|q)}{\pi_{{\theta}_{\text{old}}}(s_{i}|q)}A_{i},\text{clip}(\frac{\pi_{\theta}(s_{i}|q)}{\pi_{{\theta}_{\text{old}}}(s_{i}|q)},1-\varepsilon_{\text{low}},1+\varepsilon_{\text{high}})A_{i}\big)\Big)+\alpha\mathcal{L}_{\text{NLL}}(\theta),\end{split}$ (4)
where $G$ denotes the group size, and $A_{i}$ is the advantage of response $s_{i}$, defined as $A_{i}=\frac{\mathcal{R}_{i}-\text{mean}(\mathcal{R})}{\text{std}(\mathcal{R})}$. The reward is given by $\mathcal{R}_{i}=1$ if answer is correct, and 0 otherwise. The indicator function $I(s_{i,t})=1$ indicates that the token $s_{i,t}$ is generated by $\pi_{\theta}$ rather than observed from an external executor. Additionally, to better exploit successful trajectories, we incorporate a language modeling loss $\mathcal{L}_{\text{NLL}}$, weighted by a coefficient $\alpha$, applied to examples with positive advantage $\mathcal{T}_{\text{pos}}$, directly reinforcing the likelihood of correct samples during RL training (yue2025vapo).
$\mathcal{L}_{\text{NLL}}(\theta)=-\frac{1}{\sum_{s_{i}\in\mathcal{T_{\text{pos}}}}|s_{i}|}\sum_{s_{i}\in\mathcal{T}_{\text{pos}}}\log\pi_{\theta}(s_{i}|q).$ (5)
To stabilize training, we filter out trajectories with failed code executions during episode-level optimization to avoid inappropriate penalties. Since execution failures may arise not only from model-generated errors but also from environment issues or sandbox limitations.
Step-level Optimization. After sampling full trajectories, we perform step-level optimization to correct code errors using execution feedback, as illustrated in Figure 3(b). This stage focuses specifically on reasoning steps that resulted in failed actions $\mathcal{A}_{\text{err}}$. First, we construct a step-level optimization dataset $\mathcal{D}_{step}$. For each failed step, we treat it as a query and generate group rollouts, where the execution correctness of the generated code provides the reward signal. Consequently, it is crucial to ensure that the rollouts within each group cover diverse execution outcomes. Next, we describe the construction of $\mathcal{D}_{step}$ and its corresponding optimization method.
Within a think-act-observe tuple $(r^{t},a^{t},o^{t})$, a failed execution inherently indicates an error exists in the action $a^{t}$. Moreover, since $a^{t}$ is conditioned on $r^{t}$, fixing $r^{t}$ limits the diversity of generated actions. To overcome this problem, we implement a backtracking procedure. We partition the erroneous thought $r^{t}$ into a prefix $r^{t}_{\text{pre}}$ and a suffix $r^{t}_{\text{suf}}$ of length $L_{\text{suf}}$. Subsequently, we condition the model on the history up to $r^{t}_{\text{pre}}$ and regenerate a new reasoning suffix and action. This procedure produces the dataset $\mathcal{D}_{step}$ for fine-grained step-level optimization:
$\mathcal{D}_{step}=\{\operatorname{pref}(\tau,t)\mid a^{t}\in\mathcal{A}_{\text{err}},\tau\in\mathcal{T}\},$ (6)
$\operatorname{pref}(\tau,t)=(q,r^{1},a^{1},o^{1},...,r^{t-1},a^{t-1},o^{t-1},r^{t}_{\text{pre}}),r^{t}=r^{t}_{\text{pre}}\oplus r^{t}_{\text{suf}},$ (7)
This formulation defines a fine-grained code generation task: given a reasoning prefix $\operatorname{pref}(\tau,t)$, the model must correctly regenerate the subsequent reasoning suffix $\hat{r}^{t}_{\text{suf}}$ and action $\hat{a}^{t}$. At this stage, each sample contains a single think-act-observe loop, and the reward is computed from execution correctness. We optimize the policy with the following step-level loss:
$\begin{split}\mathcal{L}^{\text{step}}_{\pi_{\theta}}(\theta)=\mathbb{E}[&p\sim\mathcal{D}_{step},\{s_{i}^{\prime}\}_{i=1}^{G}\sim\pi_{\theta}(S|p)]\frac{1}{G}\sum_{i=1}^{G}\Big(\frac{1}{\sum_{t=1}^{|s^{\prime}_{i}|}I(s^{\prime}_{i,t})}\sum_{t:I(s^{\prime}_{i,t})=1}^{|s^{\prime}_{i}|}\\ &\min\big(\frac{\pi_{\theta}(s_{i}^{\prime}|p)}{\pi_{{\theta}_{\text{old}}}(s_{i}^{\prime}|p)}A_{i}^{\prime},\text{clip}(\frac{\pi_{\theta}(s_{i}^{\prime}|p)}{\pi_{{\theta}_{\text{old}}}(s_{i}^{\prime}|p)},1-\varepsilon_{\text{low}},1+\varepsilon_{\text{high}})A_{i}^{\prime}\big)\Big)+\alpha\mathcal{L}_{\text{NLL}}(\theta).\end{split}$ (8)
where $s^{\prime}_{i}$ and $A^{\prime}_{i}$ denote the responses and advantages at the step-level.
The final training objective combines both episode-level and step-level losses:
$\mathcal{L}(\theta)=\mathcal{L}^{\text{epis}}_{\pi_{\theta}}(\theta)+\mathcal{L}^{\text{step}}_{\pi_{\theta}}(\theta).$ (9)
2.4 Self-Correction During Inference
During inference, our model follows the think-act-observe loop. To exploit immediate feedback from tool execution, we introduce a self-correction mechanism that dynamically corrects erroneous reasoning steps, as shown in Figure 3(c). Specifically, when an action $a^{t}$ fails to execute, it indicates that both the action $a^{t}$ and its associated reasoning step $r^{t}$ are likely incorrect. To explore alternative solving paths, the model backtracks to $r^{t}$ and partitions it into a prefix $r^{t}_{\text{pre}}$ and a suffix $r^{t}_{\text{suf}}$, as previously described. Conditioned on the history up to $r^{t}_{\text{pre}}$, the model then regenerates a new reasoning suffix $\hat{r}^{t}_{\text{suf}}$ and a revised action $\hat{a}^{t}$, thereby enabling online error correction during inference. The correction procedure can be repeated for up to $N_{\text{corr}}$ attempts. Importantly, each attempt only requires regenerating the suffix $\hat{r}^{t}_{\text{suf}}$ and the corresponding action $\hat{a}^{t}$, rather than recomputing the entire trajectory. Thus, the additional computational cost is minimal compared to the total cost.
3 Experiments
Table 1: Comparison with state-of-the-art methods on mathematical benchmarks, the best results are in bold and the second-best are underlined. Code use indicates whether code tools are employed. †  denotes our reimplementation results of Avg@4, ‡  indicates results from their official releases.